where did i make my mistake in solving this problem:
x+2/x = x+2/x = 1+2=3
when i replace the x with 1 it checks out...can you please explain my error... thank you.

Answer by nerdybill(7384)

assuming you started with:
x+2/x = 3
we multiply both sides by x:
x^2+2 = 3x
x^2-3x+2 = 0
factoring the left side:
(x-2)(x-1) = 0
x = {1, 2}